OpenStudy (anonymous):

a-(9b-13b) a=12,b=-2 can someone show me how to do this plz!

OpenStudy (anonymous):

you would just plug in 12 for a and -2 for b so it wouldn't be 12-(9*-2-13*-2) and then it is just PEMDAS

OpenStudy (anonymous):

so it would be 12-(-18-26)= 12--44 or 12+44=56 (2 negatives make a positive) I hope this helped!!

OpenStudy (anonymous):

yes, thankyou

OpenStudy (anonymous):

no problem :)